OpenStudy (anonymous):

sqrt-64/(7-6i)-(2-2i) and sqrt-10(11+sqrt-5, operations of complex numbers

OpenStudy (anonymous):

anybody plz

OpenStudy (anonymous):

2 problem?

OpenStudy (anonymous):

\[\frac{-62^2}{7-6i}-(2-2i)\]?

OpenStudy (anonymous):

sorr 64 not 62

OpenStudy (anonymous):

these are two seperate equations not together, sorry if i confused you

OpenStudy (anonymous):

click on the bottom rewrite the equation

OpenStudy (anonymous):

1# sqrt-64/(7-6i)-(2-2i) 2# sqrt-10(11+sqrt-5, ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{-64}\div(7-6i)-(2-2i)\]

OpenStudy (anonymous):

\[\sqrt{-10}(11+\sqrt{-5}\]

OpenStudy (anonymous):

start second:\[\sqrt{-10}=\sqrt{-1}*\sqrt{10}\]\[\sqrt{-1}=i\] \[\sqrt{-10}*(11+\sqrt{-5)}\]\[(\sqrt{-1}*\sqrt{10}*11)+(\sqrt{-1}*\sqrt{10}*\sqrt{-1}*\sqrt{5})\]\[(11i \sqrt{10})+(i*i*\sqrt{10}*\sqrt{5})\]\[i*i=i^2=-1\]\[(11i \sqrt{10})+(-1\sqrt{50})\]\[11i \sqrt{10}-5\sqrt{2}\]